OpenStudy (anonymous):

some doctors use the formula ND= 1.08T to relate the variables N (the number of patient appointment the doctor schedules in one day), D ( the duration of each patient appointment) and T ( the total of mintues the doctors can use to see a patients in one day) (a) solve the formula for N (b) use this to find the number of patient appointment N a doctor should make if she has 6 hours available for patients and each appointment is 15 mintue long. ( round your answer to the nearest whole number)

OpenStudy (anonymous):

a) N = 1.08T/D

OpenStudy (anonymous):

b) T =6 D =15 min = 1/4 hr N = 1.08* 6 / .25

OpenStudy (anonymous):

ok for the N i need to multiply?